OpenStudy (anonymous):

Find the vertex, value of p, axis of symmetry, focus, and directrix of the parabola. y-4=-1/2(x-8)^2. Help Please.

OpenStudy (anonymous):

y - 4 = -1/2(x - 8)^2 y = -1/2(x - 8)² + 4 Vertex (8,4) <=x value opposite of inside parentheses and outside number for y a = -1/2 a = 1/(4p) -1. . . 1 ---- = ----- 2 . . .4p -4p = 2 p = -1/2 Focus (8,3.5) <=vertex plus p value onto y Directrix: y = 4.5 vertex y value - p which is - (-.5) or + .5 Axis of Symmetry: x = 8 <= x value of vertex I hope that helps!! :-)

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

Welcome!